OpenStudy (anonymous):

Can someone show me the steps on how to get this answer? I am solving nonlinear equations by substitution. y=x^2-4 y=-x^2+5 Answer is: (3times square root of 2/2, 1/2) and 3times square root of 2, 1/2)

OpenStudy (abb0t):

rearrange one of your equations so that you get both variables on one side, then substitute the other into it. Solve in terms of "x" or "y".

OpenStudy (abb0t):

\(y-x^2 = -4\) and \(y=-x^2+5\) substitue thhis equation into the top one. \((-x^2+5)-x^2=-4\) solve for "x". then sibstitue that vlue int he other.

OpenStudy (anonymous):

Can you show me the steps once you substitute? I;m getting lost in the square root part

OpenStudy (abb0t):

\(-x^2-x^2+5=-4\)

OpenStudy (anonymous):

I'm not getting how to solve it after that step. 2x^2=-9

OpenStudy (abb0t):

divide by 2 and square both sides.